OpenStudy (anonymous):

I wrote the question wrong 12^(y-2)=20

OpenStudy (solomonzelman):

1) divide both sides by 12 2) add 2 to both sides

OpenStudy (anonymous):

i wrote it wrong 12^(y-2)=20

OpenStudy (anonymous):

simplify 12y-24=20 20+24=44/12=3.67 now do reverse to check 12(3.67-2)=20 12*3.67=44-24=20

OpenStudy (anonymous):

Divide both sides by 12 then add to to both sides! Theres your answer

OpenStudy (anonymous):

Bravo noneyabusiness!

OpenStudy (anonymous):

Thanks